Given the following information for Watson Power Co., find the WACC. Assume the company’s tax rate is 21 percent. Debt: 50,000 bonds with a 4.8 percent coupon outstanding, $1,000 par value, 15 years to maturity, selling for 105 percent of par; the bonds make semiannual payments.

Common stock: 825,000 shares outstanding, selling for $72 per share; the beta is .99.

Preferred 65,000 shares of 3.8 percent preferred stock outstanding, currently stock: selling for $60 per share. Assume par value is $95.

Market: 8 percent market risk premium and 3.2 percent risk-free rate.

Can you please do this in Excel. Thank You :)
Answer #1

WACC (Weighted Average Cost of Capital)=( Total Equity/ value of the firm)* cost of Equity +(Total Debt/ Value)* cost of debt* (1- tax)+ Total Pref. share/ Value of Firm

OR = E/V*Re + D/V*Rd* (1-Tc) + P/V*RP

Value of firm= E+D+P

Given in the question

Tax rate= 21%

Debts= 50000 bonds* $1000= $50,000,000.00 (with maturity of 15 years)

Market Value of Debt= 50000 bonds* $1050 (105% of Par)= $52,500,000.00

Rd= 4.80%

Comman Stock= 825000* $72 = $59,400,000.00

Beta= 0.99

Pref. stock= 65000* $60= $3900000.00 (par value $95)

RP (rate of pref. sh)=3.80%

risk premium (Rp)= (Market rate (Rm)- risk free rate (Rf)) =8%,

Rf= 3.2%

Calculation: Re (By CAPM Model) = Rf + Risk Premium*beta= 3.2% + 8*99= 3.20% + 7.92%= 11.12%

Hence Re= 11.12% (By CAPM Model)

Value of firm= E+D+P=

$115,800,000.00

WACC= 7.5512% (Refer Excel calculation below as per formula mentioned= E/V*Re + D/V*Rd* (1-Tc) + P/V*RP)

CAPM Model
Paticulars rate
Rp 8%
Rf 3.20%
Beta 0.99
Re 11.12%
WACC
Re 11.12% Market Value of Equity $59,400,000.00
Rd 4.80% Market Value of Debt $52,500,000.00
E/V 0.512953 Market Value of Pref. share $3,900,000.00
D/V 0.453368 Value of the Firm $115,800,000.00
P/V 0.033679
Tax 21%
1-tax 79%
RP (rate of pref. sh) 3.80%
WACC 7.5512%